The length of a rectangle is 4 more than twice its width and its perimeter is 70 m. What are the dimensions of the rectangle?
Answer by macston(5194)

Width =W, Length =L=2W+4, P=70m
P=2(L+W)=2(2W+4+W)=70 m
4W+8+2W=70 m
6W+8=70 m Subtract 8 from each side
6W=62
W=10.33 Width is 10.33 m
L=2W+4=20.66+4=24.66 Length 24.66 m
CHECK
P=2(L+W)= 70 m
2(10.33+24.66)=70 m
2(34.99)=70 m
70m=70m
L=2W+4
24.66=2(10.33)+4
24.66=20.66+4
24.66=24.66